OpenStudy (anonymous):

evaluate 1/2a^-4b^-2 where a = -2 and b =4

OpenStudy (anonymous):

(-2)^-4 = 1 / 2^4 = 1/16 (4)^-2 = 1 / 4^2 = 1/16 so denominator = 2(1/16*1/16 ) = 1/128 1/1/128 = 128

OpenStudy (anonymous):

Lol answer choices 2 1/2 3 4

OpenStudy (anonymous):

problem arises because your expression can be interpreted in various ways i assumed itwas 1 / (2a^-4b^-2) have you got the original question?

OpenStudy (anonymous):

yes i do evaluate \[1/2a ^{-4}b ^{2}\] where a = -2 and b = 4

OpenStudy (anonymous):

if the 1/2 only refers to the a^-4 then we get 1/2 * 1/16 * 16 = 1/2

OpenStudy (anonymous):

Thats what i thought thanks alot :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Do you think you can look at my past questions and try the one right below it ? im pretty sure the answer is \[mq ^{2}/n ^{4}\]